# Understanding the Radial Acceleration Relation of Dwarf Galaxies with Emergent Gravity

### v1

@@ 2.2 DETERMINATION OF OBSERVED GRAVITATIONAL ACCELERATION
If we use this equation to define the mass within a characteristic radius, (e.g., the deprojected half-light radius, ) by assuming at , it becomes
where is the line-of-sight velocity dispersion111We note that total velocity dispersion, , can be expressed as line-of-sight velocity dispersion, ; see Sec.2.1 and 3.2 in Wolf et al. (2010).
Furthermore, we propose that Equation (3) be extended to any radius. As reported in Wolf et al. (2010), we can derive the relation of mass estimates between the cases of and using the relation between and :
where . Wolf et al. (2010) suggest that there could exist a radius where goes to zero (i.e. ). This means that the mass estimate is minimally affected by at this radius. On the other hand, if is zero at any radius (i.e. isotropic velocity anisotropy), the right-hand side of Equation (4) is zero. This means that we do not need to find the solution for where becomes zero. This further indicates that the mass estimates would be insensitive to at any radius; . The left-hand side of Equation (2) is independent of regardless of the radius. We also take the luminosity-weighted average on Equation (2) together with scalar virial theorem to exploit the condition (for more detail, see Wolf et al. (2010)). We estimate the observed gravitational acceleration using a simple form of the mass estimator with independence of :
We note that the observed profiles are projected, and the projected 2D radius can be replaced by the deprojected 3D radius under the assumption of spherical symmetry. As mentioned above, three galaxies in our sample (NGC 147, NGC 185, and NGC 205) are known to have non-negligible rotation (Geha et al., 2006, 2010). To take into account this rotational contribution to the gravitational acceleration, we use the following equation for such systems:
@@ 2.3 PREDICTION OF THE GRAVITATIONAL ACCELERATION VIA BARYONIC COMPONENTS
where is the gravitational acceleration from baryonic matter in the Newtonian framework as shown above, and comes from apparent dark matter, as first suggested in Verlinde (2017). We follow the method in Yoon et al. (2023) to compute the gravitational acceleration predicted from the emergent gravity. Specifically, we use the following equation to derive :
where is the characteristic acceleration scale suggested by Verlinde.
For more accurate calculation of the gravitational acceleration from the emergent gravity, we also consider the quasi-de Sitter universe as in Yoon et al. (2023) and Diez-Tejedor et al. (2018). The quasi-de Sitter universe is based on the assumptions that the positive cosmological constant represents dark energy, and that cosmic evolution explained by general relativity adequately describes the quasi-de Sitter universe. Here the quasi-de Sitter value of , , is obtained by assuming that only the cosmological constant part contributes to , which is estimated from Planck observations (Diez-Tejedor et al., 2018).
@@ Acknowledgements.
We thank Joshua D. Simon for offering the stellar-kinematic data. We thank the SNU ExGalCos team members for valuable comments. We acknowledge the support by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT), [NRF-2021R1A2C1094577(SH, HSH), NRF-2022R1A2C1092306(YY)] and by Samsung Electronic Co., Ltd. (Project Number IO220811-01945-01).References

### v2

@@ 2.2 DETERMINATION OF OBSERVED GRAVITATIONAL ACCELERATION
One can express the relation of mass estimates between the cases of and as follows (see equation (20) of Wolf et al. 2010):
where . Wolf et al. (2010) suggest that there could exist the characteristic radius, , where goes to zero. Thus, the enclosed mass is insensitive to and the value of is close to the deprojected half-light radius, . Then the mass can be expressed only with the observed line-of-sight velocity dispersion, 111The total velocity dispersion, , can be expressed with line-of-sight velocity dispersion, ; see Sec.2.1 and 3.2 in Wolf et al. (2010)., once we adopt the luminosity-weighted average on equation (2) together with scalar virial theorem to exploit the condition (see Wolf et al. 2010 for more details).
We would like to use equation (4) for any radius by assuming an isotropic velocity dispersion profile (i.e. ) to make our computation of acceleration simple. Actually there are several studies indicating that is close to zero and the model with constant is consistent with observations (Łokas, 2009); this includes the observations of dwarf spheroidal galaxies in Mashchenko (2015, Carina, Fornax, Leo I, Sculptor, and Sextan) and in Vitral et al. (2024, Draco). We can then estimate the gravitational acceleration from the observed velocity dispersion using a simple form of the mass estimator without :
Here, the third term of is from the luminosity-weighted average on equation (2) and the scalar virial theorem, and the fourth term of is from the assumption of constant and spherical symmetry. We note that the observed profiles are projected ones, and the projected 2D radius can be replaced by the deprojected 3D radius under the assumption of spherical symmetry. As mentioned above, three galaxies in our sample (NGC 147, NGC 185, and NGC 205) are known to have non-negligible rotation (Geha et al., 2006, 2010). To take into account this rotational contribution to the gravitational acceleration, we use the following equation for such systems:
@@ 2.3 PREDICTION OF THE GRAVITATIONAL ACCELERATION VIA BARYONIC COMPONENTS
where is the gravitational acceleration from baryonic matter in the Newtonian framework as shown above, and comes from apparent dark matter, as first suggested in Verlinde (2017). The above equation is derived by expanding the derivation of in Verlinde (2017), which used the theory of elastic inclusions. We follow the method in Yoon et al. (2023) to compute the gravitational acceleration predicted from the emergent gravity. Specifically, we use the following equation to derive :
where is the characteristic acceleration scale suggested by Verlinde. Note that the non-linearity manifests in the above equations. As in MOND (see Equation (12)), this violation of the superposition principle is crucial to explain the Tully-Fisher relation. Also, , which is related to the space-space component of metric, must not be confused with the Newtonian potential , related to the time-time component of metric. In the presence of spherical symmetry, the former reduces to , while the latter reduces to .
For more accurate calculation of the gravitational acceleration from the emergent gravity, we also consider the quasi-de Sitter universe as in Yoon et al. (2023) and Diez-Tejedor et al. (2018). The quasi-de Sitter universe is based on the assumptions that the positive cosmological constant represents dark energy, and that cosmic evolution explained by general relativity adequately describes the quasi-de Sitter universe. Here the quasi-de Sitter value of , , is obtained by assuming that only the cosmological constant part contributes to , which is estimated from Planck observations (Diez-Tejedor et al., 2018). However, the precise value of remains unknown, as the relativistic version of Verlinde’s emergent gravity, which is necessary to apply it to the universe as a whole in the FRW framework, has not been formulated yet. At this point, it is also unclear whether depends on time (i.e., redshift).
@@ 4 DISCUSSION
To repeat, MOND, i.e., Equation (12) underestimates the gravitational accelerations of dwarf spheroidal galaxies, in which case the emergent gravity prediction is more accurate. Therefore, considering that the gravitational accelerations of dwarf spheroidal galaxies are tiny, one may be tempted to conclude that MOND underestimates gravitational accelerations in a very low acceleration regime. However, according to the weak lensing analysis of Brouwer et al. (2021), MOND and Verlinde’s emergent gravity both predict the gravitational accelerations more or less correctly even in a very low acceleration regime. We also remark that it is well-known that MOND also underestimates the gravitational accelerations in galaxy clusters (Pointecouteau & Silk, 2005; Sanders, 2003, 2007; The & White, 1988). Interestingly enough, galaxy clusters, as well as dwarf spheroidal galaxies, are cases where the spherical symmetry is present for the distribution of matter. Thus, it is likely that MOND fails, but Verlinde’s emergent gravity succeeds in the presence of spherical symmetry. We confirmed this for spheroidal galaxies in this study, but not for galaxy clusters. Several studies have utilized Verlinde’s emergent gravity in analyzing galaxy clusters (Milgrom & Sanders, 2016; Tamosiunas et al., 2019; Islam & Dutta, 2019; Govind & Desai, 2024). However, it is uncommon to find studies that simultaneously apply both this theory and MOND to ascertain which fits more accurately with galaxy clusters. This area appears to be suitable for future investigation.
@@ Acknowledgements.
We thank the referee for helpful comments that improved the manuscript. We also thank D. Simon for offering the stellar-kinematic data and the SNU ExGalCos team members for valuable comments. We acknowledge the support of the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT), [NRF-2021R1A2C1094577(SH, HSH), NRF-2022R1A2C1092306(YY)], Samsung Electronic Co., Ltd. (Project Number IO220811-01945-01, HSH), and Hyunsong Educational & Cultural Foundation (HSH).References
@@
- Govind & Desai (2024) Govind, A., & Desai, S. 2024, J. Cosmology Astropart. Phys, 2024, 030
@@
- Łokas (2009) Łokas, E. L. 2009, MNRAS, 394, L102
- Mashchenko (2015) Mashchenko, S. 2015, arXiv e-prints, arXiv:1504.08273
@@
- Milgrom & Sanders (2016) Milgrom, M., & Sanders, R. H. 2016, arXiv e-prints, arXiv:1612.09582
@@
- Sanders (2003) Sanders, R. H. 2003, MNRAS, 342, 901
- Sanders (2007) Sanders, R. H. 2007, MNRAS, 380, 331
@@
- Tamosiunas et al. (2019) Tamosiunas, A., Bacon, D., Koyama, K., & Nichol, R. C. 2019, J. Cosmology Astropart. Phys, 2019, 053
- The & White (1988) The, L. S., & White, S. D. M. 1988, AJ, 95, 1642